# Vulnerabilities of smart contracts and mitigation schemes: A Comprehensive Survey

### v1

@@ V-E1 Description
Echidna is a specialised property-based fuzzing tool for smart contracts implemented by Grieco at al. [9] that takes inspiration from QuickCheck [94]. Echidna seeks to violate user-defined invariants that represent potential contract errors rather than just finding crashes [95]. These invariants are writtern by the tester and it is called echidna property: A special Solidity function with no arguments that returns ”true” on success and has a name that begins with ”echidna”. Any transactions that result in these properties returning false or error are reported by Echidna, essentially disclosing contract bugs [96]. Echidna employs a variety of strategies, including iterative feedback, structural constraints, and known input variation, to produce test inputs.
V-E2 Execution Example
@@
In order to test Echidna, we have adapted the tokensalechallenge.sol smart contract [85]. We created a derived contract TestTokenSaleChallenge that extends the original TestTokenSaleChallenge contract shown in Listing 22.The test properties in the TestTokenSaleChallenge are designed to check for arithmetic errors that can occur in the logic used for NFT fractionalization.
@@
Echidna generates its results directly to the console, as figure 23 shows, distinguishing between properties that were successfully validated and those that failed. For each property that didn’t meet the validation criteria, Echidna offers a counterexample, which is essentially a step-by-step breakdown of how the property failed under specific conditions. In our example, the underflow test did not pass, even without initiating a transaction. This was because the msg.sender was left empty, indicating that the system can encounter failures in scenarios where sender information is not provided. However Echidna didn’t detect an overflow issue, possibly due to a short timeout period, preventing it from producing a sequence revealing this vulnerability as we did not change the default configuration.
V-E3 Pros and Cons
@@
One of Echidna’s notable strengths lies in its straightforward setup process, offering users multiple installation options, including direct build with Stack, Docker, or Homebrew. This flexibility facilitates accessibility across different platforms and development environments.
@@
However, users are required to explicitly define the properties they wish to test. This feature allows for customized and focused testing scenarios, ensuring that the analysis is directly relevant to the contract’s intended behaviors and security assumptions. Despite the advantage of tailored testing, this approach adds a layer of complexity. Users must possess a thorough understanding of their contract’s logic and potential vulnerabilities to effectively define meaningful test properties. This necessity for manual property definition demands a higher level of engagement and expertise from the user, compared to tools that automatically infer test cases or vulnerabilities.
V-F Summary
@@
Table VII presents a comparative analysis of Oyente, Slither, Mythril, Manticore and Echidna. The comparison is based on the number of detected vulnerabilities, used methodology, an academic or company tool, the code level, required solidity version, and availability of documentation.
Oyente | Slither | Mythril | Manticore | Echidna | |
Number of detected vul | 4 | 70 | 14 | 10 | - |
Methodology | Symbolic | Symbolic | Static | Symbolic | Fuzzing |
Code Level | EVM | Solidity | Solidity | EVM | Solidity |
Restrictions (Solidity) | 0.4 | 0.4 | 0.4 | 0.4 | not specified |
Documentation | Low | Medium | High | High | Medium |
@@
In conclusion, there are a variety of tools that can be used to test smart contracts and identify a wide range of vulnerabilities. Static analysis tools are still the most popular because they are simple to use compared to (a) fuzzing which necessitates a lot of resources to set up the test environment; and (b) formal formal verification which necessitates expertise in contract specification writing.
VI Guidelines for secure smart contracts
@@
Writing a secure and bug-free smart contract is crucial as a small bug can lead to the loss of millions of dollars. In this section, we provide guidelines on how to write a secure smart contract [97]. These steps are summarized in Figure 24
@@
The first step to test a smart contract is to visualize its control flow right in order to have a global view and better understanding of the interactions between its different components. Slither is an interesting tool to use during this step since it has a variety of printers that can be used to visualise call graphs, contract inheritance relationships, modifiers called by each function, etc. Solidity Visual Developer is another interesting tool that could be integrated into Visual Studio Code [98]. This extension helps with semantic highlighting. It also generates a detailed class outline. This step is important to spot critical functions that must be tested in the next stage.
@@
Afterwards, the smart contract should be subjected to an automatic analysis. It is best to start with common bug detection using Mythril and Slither, we recommend combining tools for better results. To target a certain critical function, it is best to use Manticore tool as it allows the implementation of specific use cases. It is also crucial to employ dynamic analysis tools to decrease the number of false positive generated by static tools and to cover additional states. Echidna is an interesting ethereum fuzzing tool to use [9]. It is recommended to look for unique features in smart contracts. For instance, if the smart contract is an ERC token, it will be important to use Slither’s ERC Conformance functionality to make sure that the smart contract conforms to the ERC standards put in place.
@@
Because automated tools are unaware of the context for which the contract was produced, running more specific testing will increase the detection accuracy. To achieve this, Echidna and Manticore [99] allow for the definition of security properties in solidity. Indeed, these tools are useful for checking arithmetic operations, external interactions, and standardisation. Whereas Slither allows the definition of properties with Slither Python API’s, it allows checking for inheritance, variable dependencies, and access restrictions.
@@
The third step will be formal verification; it is crucial since it ensures that the contract requirements are legitimate. This step, however, is challenging as it requires special expertise (e.g., writing the formal specification using a specification language). Finally, the output of the preceding steps must be manually verified and corrective actions based on patterns and mitigation techniques must be applied. After fixing the bugs, one has to redo the previous steps to increase the probability of a bug free contract.
VII Related Work
@@
In the last few years, several of surveys covering smart contract vulnerabilities have been published. The majority of these surveys cover one or two of the following areas: Security vulnerabilities of smart contracts, analysis of existing tools and methodologies to detect vulnerabilities of smart contracts, formal specification and verification, and common design patterns and practices of smart contracts. Kushwaha et al.[100] conducted a systematic review of Ethereum smart contract analysis tools, classifying them into static and dynamic analysis categories. They scrutinize 86 tools, discussing their methodologies, such as taint analysis, symbolic execution, and fuzzing. However, it lacks of consideration for the developer community’s preferences and practical scenarios where each analysis tool would be most beneficial. Kushwaha et al. [21] presented a comprehensive systematic review of security vulnerabilities in Ethereum blockchain smart contracts. The paper categorizes vulnerabilities into three main root causes and seventeen sub-causes, providing in-depth insights into twenty-four specific vulnerabilities along with their prevention methods, detection, and analysis tools. Despite providing a comparative analysis of various detection tools, the paper falls short in offering in-depth feedback on the effectiveness and limitations of these tools. Di Angelo et al. [101] provided a state-of-the-art review of analysis tools of Ethereum smart contracts. The study is based on real execution of the tools. It classifies the tools based on availability, maturity level, purpose, and analysis method. However, the survey [101] become obsolete (e.g., no longer maintained). Harz et al. [102] presented languages, paradigms, and a verification approach for smart contracts. They did not; however, go into depth about the verification tools or known vulnerabilities. Li et al. [103] discussed blockchain safety problems and proposed enhancements, however, the survey [103] is general because, for example, it included different types of blockchain, such as Ethereum and bitcoin. Saad et al. [104] investigated several attacks on various blockchain platforms, as well as protection mechanisms. They briefly covered mitigation schemes, however, they did not cover tools to detect vulnerabilities. Atzei et al. [24] presented major vulnerabilities and attacks related to Ethereum smart contracts, however, they did not cover mitigation schemes. Chen et al. [105] presented 40 types of Ethereum vulnerabilities, along with their causes and 29 attacks; however, they did not cover tools to detect vulnerabilities. Zhu et al. [106] presented 11 smart contract vulnerabilities in depth, along with various defences against well-known attacks. They did cover schemes to detect those vulnerabilities. Durieux et al. [107] conducted an empirical review of nine automated analysis tools on 47,518 contracts, They developed a framework to analyze these tools. The analysis makes use of two data sets of smart contracts with tagged vulnerabilities. However, the study did not cover vulnerabilities of smart contracts and mitigation schemes.
@@
VIII Conclusion
In this study, we presented detailed common vulnerabilities in Ethereum smart contracts, as well as mitigation solutions based on patterns and standards. We have covered the most popular detection methodologies, such as static analysis, dynamic analysis, and formal verification. We also did discuss the benefits and drawbacks of each methodology. In addition, we provided community-recommended vulnerability detection tools, along with an execution sample and detailed feedback for each. Based on our investigation of existing methodologies and tools, we proposed a guideline with the tool(s) to be used for each step in auditing smart contracts.
We did observe that most of the tools detect vulnerabilities but do not provide refactoring recommendations. In future work, we aim to implement a refactoring module that proposes improvements based on the detected bugs and error traces. We also aim in study attacks related to NFTs such as wash trading which is a type of market manipulation where attackers repeatedly buy and sell the same NFT to increase the price and trading volume.

### v2

@@ 5.5.1 Description
Echidna is a specialized property-based fuzzing tool for smart contracts implemented by Grieco at al. [9] that takes inspiration from QuickCheck [94]. Echidna seeks to violate user-defined invariants that represent potential contract errors rather than just finding crashes [95]. The tester writes these invariants, and it is called echidna property: A particular Solidity function with no arguments that returns ”true” on success and has a name that begins with ”echidna.” Echidna reports any transactions that result in these properties returning false or error, essentially disclosing contract bugs [96]. Echidna employs various strategies to produce test inputs, including iterative feedback, structural constraints, and known input variation.
5.5.2 Execution Example
@@
To test Echidna, we have adapted the tokensalechallenge.sol smart contract [85]. We created a derived contract TestTokenSaleChallenge that extends the original TestTokenSaleChallenge contract shown in Listing 22. The test properties in the TestTokenSaleChallenge are designed to check for arithmetic errors that can occur in the logic used for NFT fractionalization.
@@
Echidna generates its results directly to the console, as figure 23 shows, distinguishing between successfully validated and failed properties. Echidna offers a counterexample for each property that didn’t meet the validation criteria, which is essentially a step-by-step breakdown of how the property failed under specific conditions. In our example, the underflow test did not pass, even without initiating a transaction. This was because the msg.sender was left empty, indicating that the system can encounter failures in scenarios where sender information is not provided. However, Echidna didn’t detect an overflow issue, possibly due to a short timeout period, which prevented it from producing a sequence revealing this vulnerability as we did not change the default configuration.
5.5.3 Pros and Cons
@@
One of Echidna’s notable strengths is its straightforward setup process. It offers users multiple installation options, including direct build with Stack, Docker, or Homebrew. This flexibility facilitates accessibility across different platforms and development environments.
@@
However, users are required to define the properties they wish to test explicitly. This feature allows for customized and focused testing scenarios, ensuring the analysis is directly relevant to the contract’s intended behaviors and security assumptions. Despite the advantage of tailored testing, this approach adds a layer of complexity. Users must thoroughly understand their contract’s logic and potential vulnerabilities to define meaningful test properties effectively. This necessity for manual property definition demands a higher level of engagement and expertise from the user than tools that automatically infer test cases or vulnerabilities.
5.6 Summary
@@
Table VII presents a comparative analysis of Oyente, Slither, Mythril, Manticore, and Echidna. The comparison is based on the number of detected vulnerabilities, the methodology used, an academic or company tool, the code level, the required solidity version, and the availability of documentation.
Oyente |
Slither |
Mythril |
Manticore |
Echidna |
Number of detected vul |
4 |
70 |
14 |
10 |
Methodology |
Symbolic |
Symbolic |
Static |
Symbolic |
Fuzzing |
Code Level |
EVM |
Solidity |
Solidity |
EVM |
Solidity |
Restrictions (Solidity) |
0.4 |
0.4 |
0.4 |
0.4 |
not specified |
Documentation |
Low |
Medium |
High |
High |
Medium |
@@
In conclusion, there are various tools that can be used to test smart contracts and identify a wide range of vulnerabilities. Static analysis tools are still the most popular because they are simple to use compared to (a) fuzzing, which necessitates a lot of resources to set up the test environment, and (b) formal verification, which necessitates expertise in contract specification writing.
6 Guidelines for secure smart contracts
@@
Writing a secure and bug-free smart contract is crucial, as a small bug can lead to the loss of millions of dollars. In this section, we provide guidelines on how to write a secure smart contract [97]. These steps are summarized in Figure 24
@@
The first step in testing a smart contract is to visualize its control flow correctly to have a global view and better understand the interactions between its different components. Slither is an exciting tool to use during this step since it has a variety of printers that can be used to visualize call graphs, contract inheritance relationships, modifiers called by each function, etc. Solidity Visual Developer is another interesting tool that could be integrated into Visual Studio Code [98]. This extension helps with semantic highlighting. It also generates a detailed class outline. This step is essential to spot critical functions that must be tested in the next stage.
@@
Afterward, the smart contract should be subjected to automatic analysis. It is best to start with common bug detection using Mythril and Slither; we recommend combining tools for better results. To target a certain critical function, it is best to use the Manticore tool as it allows the implementation of specific use cases. It is also crucial to employ dynamic analysis tools to decrease the number of false positives generated by static tools and to cover additional states. Echidna is an interesting ethereum fuzzing tool to use [9]. It is recommended to look for unique features in smart contracts. For instance, if the smart contract is an ERC token, it will be essential to use Slither’s ERC Conformance functionality to make sure that the smart contract conforms to the ERC standards put in place.
@@
Because automated tools are unaware of the context for which the contract was produced, running more specific testing will increase the detection accuracy. To achieve this, Echidna and Manticore [99] allow for the definition of security properties in solidity. Indeed, these tools help check arithmetic operations, external interactions, and standardization. Whereas Slither allows the definition of properties with Slither Python APIs, it will enable checking for inheritance, variable dependencies, and access restrictions.
@@
The third step will be formal verification, which ensures that the contract requirements are legitimate. This step, however, is challenging as it requires unique expertise (e.g., writing the formal specification using a specification language). Finally, the output of the preceding steps must be manually verified, and corrective actions based on patterns and mitigation techniques must be applied. After fixing the bugs, one has to redo the previous steps to increase the probability of a bug-free contract.
7 Related Work
@@
Several surveys covering smart contract vulnerabilities have been published in the last few years. The majority of these surveys cover one or two of the following areas: smart contract security vulnerabilities, analysis of existing tools and methodologies to detect smart contract vulnerabilities, formal specification and verification, and common design patterns and practices. Kushwaha et al.[100] conducted a systematic review of Ethereum smart contract analysis tools, classifying them into static and dynamic analysis categories. They scrutinize 86 tools, discussing their methodologies, such as taint analysis, symbolic execution, and fuzzing. However, it lacks consideration for the developer community’s preferences and practical scenarios where each analysis tool would be most beneficial. Kushwaha et al. [21] presented a comprehensive systematic review of security vulnerabilities in Ethereum blockchain smart contracts. The paper categorizes vulnerabilities into three main root causes and seventeen sub-causes, providing in-depth insights into twenty-four specific vulnerabilities and their prevention methods, detection, and analysis tools. Despite providing a comparative analysis of various detection tools, the paper falls short in offering in-depth feedback on the effectiveness and limitations of these tools. Di Angelo et al. [101] provided a state-of-the-art review of analysis tools of Ethereum smart contracts. The study is based on the actual execution of the tools. It classifies the tools based on availability, maturity level, purpose, and analysis method. However, the survey [101] becomes obsolete (e.g., no longer maintained). Harz et al. [102] presented languages, paradigms, and a verification approach for smart contracts. They did not, however, go into depth about the verification tools or known vulnerabilities. Li et al. [103] discussed blockchain safety problems and proposed enhancements; however, the survey [103] is general because, for example, it included different types of blockchain, such as Ethereum and Bitcoin. Saad et al. [104] investigated several attacks on various blockchain platforms and protection mechanisms. They briefly covered mitigation schemes. However, they did not cover tools to detect vulnerabilities. Atzei et al. [24] presented significant vulnerabilities and attacks related to Ethereum smart contracts. However, they did not cover mitigation schemes. Chen et al. [105] presented 40 types of Ethereum vulnerabilities, their causes, and 29 attacks; however, they did not cover tools to detect vulnerabilities. Zhu et al. [106] presented 11 smart contract vulnerabilities in-depth, along with various defenses against well-known attacks. They covered schemes to detect those vulnerabilities. Durieux et al. [107] conducted an empirical review of nine automated analysis tools on 47,518 contracts; they developed a framework to analyze these tools. The analysis uses two data sets of smart contracts with tagged vulnerabilities. However, the study did not cover vulnerabilities of smart contracts and mitigation schemes.
@@
8 Conclusion
In this study, we presented detailed common vulnerabilities in Ethereum smart contracts and mitigation solutions based on patterns and standards. We have covered the most popular detection methodologies: static analysis, dynamic analysis, and formal verification. We also discussed the benefits and drawbacks of each method. In addition, we provided community-recommended vulnerability detection tools, an execution sample, and detailed feedback for each. Based on our investigation of existing methodologies and tools, we proposed a guideline with the tool(s) for each step in auditing smart contracts.
We observed that most tools detect vulnerabilities but do not provide refactoring recommendations. In future work, we aim to implement a refactoring module that proposes improvements based on the detected bugs and error traces. We also aim to study attacks related to NFTs, such as wash trading, a type of market manipulation in which attackers repeatedly buy and sell the same NFT to increase the price and trading volume.